thier are rabits and chiken in the farrm. then the farmer count all his animals and resulted to 100 both heads and feets. how many animals chicken and rabits are thier?
Answer by ankor@dixie-net.com(12701)

their are rabbits and chicken in the farm. then the farmer count all his animals and resulted to 100 both heads and feet. how many animals chicken and rabbits are their?
:
Let r = no. of rabbits
Let c = no. of chickens
:
head + feet = 100
r + 4r + c + 2c = 100
5r + 3c = 100
3c = 100 - 5r
c = (100/3) - (5/3)r
:
Choose a low integer value for r, check to see if c is an integer
There are several chicken/rabbit solutions;
2/30
5/25
8/20
11/15
14/10
17/5
20/0
: